# Enhancing the Generalization Capability of Skin Lesion Classification Models with Active Domain Adaptation Methods

### v1

@@ Abstract
We propose a method to improve the generalization ability of skin lesion classification models by combining self-supervised learning (SSL), unsupervised domain adaptation (UDA), and active domain adaptation (ADA). The main steps of the approach include selection of a SSL pretrained model on natural image datasets, subsequent SSL retraining on all available skin lesion datasets, finetuning of the model on source domain data with labels, application of UDA methods on target domain data, and lastly, implementation of ADA methods. The efficacy of the proposed approach is assessed across ten skin lesion datasets of domains, demonstrating its potential for enhancing the performance of skin lesion classification models. This approach holds promise for facilitating the widespread adoption of medical imaging models in clinical settings, thereby amplifying their impact.
1 Introduction
@@
The past decade has witnessed a surge in the application of artificial intelligence (AI) to skin cancer detection. Given the growing incidence of skin cancer within an aging population and the scarcity of dermatology experts, there is a compelling demand for automated AI solutions [60]. Globally it’s estimated that around three billion people lack adequate access to skin disease medical care [15]. Using AI can greatly alleviate the problem. Dermoscopy images are normally used by dermatologists to detect skin cancers. So computer vision (CV) methods have been applied for skin cancer classification and segmentation [54, 38, 37]. Since the start of annual challenges organized by the International Skin Imaging Collaboration (ISIC) in 2016, steady progress has been made in algorithm performance, with deep learning models surpassing human experts on the provided datasets [60, 14].
@@
Despite these advancements in medical imaging applications, model generalization remains a crucial hurdle for broad clinical adoption [35, 69, 14, 26]. For skin cancer detection specifically, Barros et al. and Daneshjou et al. find that models perform worse on dark skin tones than light skin tones, and worse on uncommon disease types [5, 16]. Han et al. observe that the dermatologist-level model’s performance significantly decreases when testing on out-of-distribution data [26]. The differences between skin lesion datasets are pervasive, stemming from various sources such as demographics, skin lesion locations, institutions, acquisition devices, and lighting conditions. Transfer learning (TL) is a technique to transfer a model’s knowledge from one domain to others. Due to the lack of large labeled datasets, a common approach is to initially train models on natural image datasets, such as imagenet [17, 47, 49]. Subsequently, the model is fine-tuned on medical image datasets. There are various factors affecting a model’s transferability, including transfer techniques, training data size, domains’ similarity, and model architecture [46]. Here we focus on the transfer techniques, since some of the factors are out of our control in actual use cases, e.g. we have to work with a given trained model on a given dataset.
@@
One transfer learning technique is using self-supervised learning (SSL). Supervised learning (SL) encourages models to learn task-specific features, potentially hindering their ability to generalize to unseen data. SSL, on the other hand, enables models to learn general features that are more robust across diverse datasets [2]. Azizi et al. propose a robust and data-efficient generalization framework for medical imaging applications by doing SSL training on in-domain medical data prior to finetuning. This approach outperforms the SL models on both in-domain and out-of-domain data. Haggerty et al. confirm the efficacy of this approach on skin lesion datasets, demonstrating that SSL retraining on skin lesion data enhances the performance of a model pretrained on imagenet data with SL or SSL methods [25]. Another TL approach is domain adaptation (DA). DA methods aim to align the source domain (data on which the model is trained) and the target domain (data the model hasn’t encountered). These are few works on applying unsupervised domain adaptation (UDA) methods for skin lesion classification [65, 9, 24].
@@
In this study, we aim to combine the SSL and DA for better transfer learning. To the best of our knowledge, we are the first to apply this approach for skin lesion classification. More specifically, we use the state-of-the-art (SOTA) SSL method, DINO, which is different from the Barlow Twins SSL method used in [2]. Besides using the UDA method, we also apply active domain adaptation (ADA) [45] methods. This is motivated by the actual clinical setting where the model can get feedback from human experts to iteratively improve performance [63]. There is little work on combining the SSL and DA methods. Zhao et al. combines SSL and AL without domain adaptation for skin lesion segmentation application [71]. Xu et al. combines SSL and UDA without ADA, and they use the SSL method of predicting rotations [67]. We demonstrate the effectiveness of the approach in the skin lesion classification application. This approach can easily integrate into the clinical workflow and iteratively improve the model performance by incorporating information from human experts. The approach consists of the four main steps with the initial preparation step of selecting a SSL pretrained model on imagenet. 1. Apply SSL retraining on all available skin lesion datasets. 2. Finetuning the model on the source domain data with labels. 3. Apply unsupervised domain adaptation on the target domain data. 4. Apply active domain adaptation as needed. This step requires input from human experts. They can decide how many labels to provide.
@@
  • We propose a workflow to iteratively improve the generalization capacity of the skin lesion classification model across domains by combining SSL training, UDA and ADA methods.
  • We do an ablation study to find the performance contribution of each step of the workflow, the SSL training, UDA and ADA.
  • Through extensive experiments, we show that our approach achieves SOTA generalization performance on ten skin lesion target domains’ data.
2 Related work
2.1 Self-supervised learning
@@
SSL is a technique to train models without labels. It can be used to create foundational models which can be finetuned for different downstream tasks. The success of SSL is initially observed in natural language processing (NLP) [18, 6], and later extended to CV [19, 20, 23, 39, 50]. Recently there has been a rapid proliferation of SSL applications in the medical imaging field [3, 41, 55, 56, 66, 32]. SSL can achieve better performance than SL trained models with much less labels [10]. There are four main types of SSL methods, innate relationships, generative, contrastive and self-prediction. Innate relationships leverage the internal structure of the data, e.g. predicting the rotation angle, or finding the position of shuffled image patches. Generative methods use autoencoders or generative adversarial networks (GANs). They are assessed by the quality of reconstructing the original images. Contrastive methods rely on using data augmentation to create positive pairs of an image. Different images are treated as negative pairs. There are various distance metrics being proposed trying to keep positive pairs close and negative pairs far away in the embedding space. Some examples include SimCLR [10], MoCo [27], BYOL [4], SimSiam [11], DINO [7]. Self-prediction methods mask out portions of the original image and try to reconstruct the original image. The main difference between self-prediction and generative methods is that self-prediction applies augmentation on portions of the original image while generative methods apply to the whole image.
SSL application in medical imaging is rapidly rising, since medical labels are costly and time consuming to get. SSL can speed up the process of model development. When applying SSL for medical imaging, special attention is needed when choosing the appropriate data augmentation methods. In general, Morningstar et al. shows that augmentation diversity contributes more to SSL performance compared to algorithmic changes [64]. However, due to the difference between natural images and medical images, Morita et al. shows that employing simple transformations including the left/right, up/down flipping, rotation and region cropping with a large ratio of the raw image can improve the classification performance by more than 3% compared with more complicated data augmentations [48].
2.2 Unsupervised domain adaptation
@@
UDA methods can be divided into two types, moment matching and adversarial training. Moment matching methods try to match the first or second moments between the hidden activation distributions of the source and target domains [70]. Some methods include deep adaptation networks (DAN) [42], joint adaptation networks (JAN) [43], correlation alignment [58], and central moment discrepancy (CMD) [70]. Adversarial training methods try to learn domain invariant features. It trains a domain classifier to put the source and target domain features in the same space. The most used adversarial training method is domain adversarial neural networks (DANN) [22]. Other methods build upon it such as adversarial discriminative domain adaptation (ADDA) [62], maximum classifier discrepancy (MCD) [52], batch spectral penalization (BSP) [12] and minimum class confusion (MCC) [34].
@@
There are works of applying UDA in medical image analysis focusing on classification [1], object localization, and segmentation [36]. More specifically for skin lesion classification, Chamarthi et al. compare performance of different UDA methods [9]. Wang et al. show applying UDA methods can mitigate model bias against minority groups [65].
2.3 Active domain adaptation
@@
UDA methods’ performance is often worse than their supervised counterpart [59, 13]. Supervised domain adaptation (SDA) requires labels of the target domain. However, for the medical imaging field especially, obtaining large amounts of labels can be cost intensive. Thus how to select the most informative samples to annotate becomes important. This is where active learning (AL) comes in. AL is trying to maximize a model’s performance given a limited amount of labels. Active domain adaption (ADA) combines active learning with domain adaptation. Different active learning methods have different sampling strategies. Broadly they can be separated into a few categories, such as uncertainty [40, 53], diversity [33, 30], representativeness [68]. There are methods trying to combine multiple types of metrics [51, 31, 21]. Zhao et al. applied SSL and AL for skin lesion segmentation, but there is no domain adaptation [71]. There is no prior work on applying ADA methods to skin lesion classification problems.
3 Methodology
@@
Here we propose the combination of SSL and DA methods into one workflow to improve skin lesion classification model’s generalization ability. It consists of the four steps.
Figure 1: Proposed workflow for model generalization. There are 4 steps, SSL retraining on in-domain data of a SSL pretrained model , finetuning the model on source domain, applying UDA, and lastly applying ADA.
@@
The preparation step is selecting a SSL pretrained model on natural image datasets, e.g. imagenet. Here we select the DINO SSL method since it’s shown to produce the SOTA performance on imagenet among many SSL methods, e.g. SimCLR, MoCov2, Barlow Twins, BYOL, SwAV [44, 7]. It uses knowledge distillation with a student and teacher architecture. The loss term minimizes the cross-entropy between the student and teacher learned features. It applies centering and sharpening on the teacher’s output to avoid learning trivial solutions. It doesn’t require large batch size, or negative samples. We choose to use the ResNet50 [28] backbone model architecture since it’s used in other similar studies [9], and we can compare results. The DINO cross-entropy loss term between the teacher and student learned features:
@@
The first step is to continue DINO training on skin lesion datasets. Since no label is needed, we can train it on all the available skin lesion data. This can help the model generalize to skin lesion data from natural images.
The second step is finetuning the model on skin lesion dataset with labels. Here we choose a very common task, binary classification of melanoma (cancer) vs. nevus (benign). We keep the backbone model frozen and only finetune the linear classification layer, because we want to directly compare the learned feature quality without the confounding factor of a more complex classifier.
The third step is applying UDA to align the source and target domain. The UDA method we choose is DANN [22]. It’s a classic UDA method and has shown good performance among UDA methods applied to skin lesion datasets [9]. DANN is an adversarial training based method. It has three main components, feature extractor, class classifier, and domain classifier. which tries to learn domain invariant features with a gradient reversal layer.
$L=min_{f},max_{d}Lc(G_{y}(G_{f}(x)),y)+Ld(G_{d}(G_{f}(x)),d)$
Lc and Ld are the cross-entropy loss of skin lesion classification and domain classification receptively. y and d are the true skin cancer label and the true domain label. $Gd,Gy,Gf$ are the networks of domain classifier, class classifier and feature extractor respectively. Their parameters are $d,y,f$ respectively.
The fourth step is applying ADA. We choose the ADA method, AADA [57] since it incorporates the DANN method. The sampling criterion is as such:
@@
It includes the diversity cue $\frac{1-G_{d}(G_{f}(x))}{G_{d}(G_{f}(x))}$ and uncertainty cue $H(G_{y}(G_{f}(x)))$. The pseduo code of applying AADA method is show below.
Algorithm 1 AADA
Input: labeled source $L_{s}=H$; unlabeled target $U_{t}\in\{HA,HLH,HLP,B,BA,BLH,BLP,M,MA,MLH\}$; labeled target $L_{t}=\emptyset$; budget per round $b=10$.
Model: =$\mathcal{M}=\{G_{f},G_{y},G_{d}\}$; feature extractor $G_{f}$; class predictor $G_{y}$; discriminator $G_{d}$. $G_{f}$ is DINO trained on skin lesion datasets and finetuned on dataset, H.
UDA: Train $\mathcal{M}$ with ($L_{s}$, $U_{t}$) using DANN loss
for $round\leftarrow 1$ to $MaxRound=10$ do
     Compute $S(x)$ $\forall x\in U_{t}$
     Select a set of $b$ images $z$ from $U_{t}$ according to $S(z)$
     Get labels $y_{z}$ from human experts
     $L_{t}\leftarrow L_{t}\cup(z,y_{z})$
     $U_{t}\leftarrow U_{t}\setminus(z,y_{z})$
     Train $\mathcal{M}$ with ($L_{s}\cup L_{t}$, $U_{t}$)
end for
4 Experiment
4.1 Dataset
[35] provide skin lesion datasets of different domains by grouping data from three datasets, HAM10000 [61], BCN20000 [29], and MSK [8], according to biological traits, such as patient age and lesion location. They quantify the intensity of domain shifts with divergence metrics, cosine similarity and JS divergence. The details of the datasets are in Table 1.
Abbreviation Origin Biological factors Melanoma amount Nevus amount Total target size
@@ 4.1 Dataset
Table 1: Skin lesion datasets of different domains. Here lists a total of 11 datasets. Dataset H is the source domain, and other datasets are target domains. It also shows the number of samples in each dataset, the melanoma and nevus ratio, and biological traits.
4.2 Metrics
We choose to use AUPRC metric because it accounts for unbalanced classes datasets well and it emphasizes more on the cancer class. As we can see from Table 1, many of the skin lesion datasets have very high levels of unbalanced classes with cancer class as low as 4%. We don’t use accuracy or AUROC, because they only work well for balanced classes datasets. However, we do need to be mindful of the different AUPRC baselines of different datasets, which are the melanoma ratios.
4.3 Model Training

### v2

# Enhancing Skin Lesion Classification Generalization with Active Domain Adaptation

@@ Abstract
We propose a method to improve the generalization of skin lesion classification models by combining self-supervised learning (SSL) and active domain adaptation (ADA). The main steps of the approach include selection of an SSL pre-trained model on natural image datasets, subsequent SSL retraining on all available skin-lesion datasets, fine-tuning of the model on source domain data with labels, and application of ADA methods on target domain data. The efficacy of the proposed approach is assessed in ten skin lesion datasets with five different ADA methods, demonstrating its potential to improve generalization in settings with different amounts of domain shifts.
Clinical Relevance—This approach is promising in facilitating the widespread clinical adoption of deep learning models for skin lesion classification, as well as other medical imaging applications.
I INTRODUCTION
@@
The past decade has witnessed a surge in the application of artificial intelligence (AI) to skin cancer detection. Given the increasing incidence of skin cancer within an aging population and the scarcity of dermatology experts, a compelling demand exists for automated AI solutions [68]. Globally, it’s estimated that around three billion people lack adequate access to skin disease medical care [16]. Using AI can greatly alleviate the problem. Dermoscopy images are normally used by dermatologists to detect skin cancers. Consequently, computer vision (CV) methods have been applied for skin cancer classification and segmentation [61, 44, 43]. Since the International Skin Imaging Collaboration (ISIC) began organizing annual challenges in 2016, steady progress has been made in algorithm performance. Deep learning models now exceed human expert performance on the provided datasets [68, 15].
@@
Despite these advances in medical imaging applications, model generalization remains a crucial hurdle for broad clinical adoption [40, 77, 15, 31, 75]. For skin cancer detection specifically, Barros et al. and Daneshjou et al. find that models perform worse on dark skin tones than light skin tones, and worse on rare diseases [6, 17]. Han et al. observe that the dermatologist-level model’s performance significantly decreases when testing on out-of-distribution data [31]. The differences between skin lesion datasets are pervasive, stemming from various sources such as demographics, skin lesion locations, institutions, acquisition devices, and lighting conditions. Transfer learning (TL) is a technique to transfer a model’s knowledge from one domain to another. Due to the lack of large labeled datasets, a common approach is to initially train models on natural image datasets, such as ImageNet [18, 53, 54]. Subsequently, the model is fine-tuned on medical image datasets. There are various factors affecting a model’s transferability, including transfer techniques, training data size, domains’ similarity, and model architecture [52]. Here we focus on the transfer techniques, since some of the factors are out of our control in actual use cases, e.g. we have to work with a given trained model on a given dataset.
@@
One transfer learning technique is using self-supervised learning (SSL). Supervised learning (SL) encourages models to learn task-specific features, potentially hindering their ability to generalize to unseen data. SSL, on the other hand, enables models to learn general features that are more robust across diverse datasets [3]. Azizi et al. propose a robust and data-efficient generalization framework for medical imaging applications by doing SSL training on in-domain medical data prior to finetuning. This approach outperforms the SL models on both in-domain and out-of-domain data. Kang et al. demonstrate similar results for four different SSL methods [42]. Haggerty et al. confirm the efficacy of this approach on skin lesion datasets, demonstrating that SSL retraining on skin lesion data enhances the performance of a model pre-trained on ImageNet data with SL or SSL methods [30]. Another TL approach is domain adaptation (DA). DA methods aim to align the source domain (data on which the model is trained) and the target domain (data the model hasn’t encountered). Few studies have explored the application of unsupervised domain adaptation (UDA) methods for skin lesion classification [73, 10, 28].
@@
In this study, we aim to combine the SSL and DA for better transfer learning. Limited research exists on combining the SSL and DA methods. Zhao et al. combine SSL and AL without domain adaptation for skin lesion segmentation application [79]. Xu et al. combines SSL and UDA for object detection and segmentation [76]. To the best of our knowledge, this work represents the first attempt to combine SSL and DA for skin lesion classification. More specifically, we use the state-of-the-art (SOTA) SSL method, DINO, which is different from the Barlow Twins SSL method used in [3]. Instead of using the UDA method, we apply active domain adaptation (ADA) [51] methods. This is motivated by the actual clinical setting where the model can get feedback from human experts to iteratively improve performance [71]. We demonstrate the effectiveness of the approach in the skin lesion classification application. This approach offers seamless integration into the clinical workflows, enabling iterative performance improvement by incorporating feedback from human experts.
@@
  • We propose a workflow to improve the generalization of the skin lesion classification model across domains by combining SSL and ADA methods.
  • We show that SSL can be an effective UDA method and applying ADA methods can add further improvements.
  • We demonstrate our approach’s efficacy on ten skin lesion target domains’ data with five different ADA methods.
II Related work
II-A Self-supervised learning
@@
SSL is a technique to train models without labels. It can be used to create foundational models which can be fine-tuned for different downstream tasks. The success of SSL is initially observed in natural language processing (NLP) [19, 7], and later extended to CV [20, 21, 27, 45, 55]. Recently there has been a rapid proliferation of SSL applications in the medical imaging field [4, 47, 63, 64, 74, 37]. SSL can achieve better performance than SL trained models with much less labels [11]. There are four main types of SSL methods, innate relationships, generative, contrastive and self-prediction. Innate relationships leverage the internal structure of the data, e.g. predicting the rotation angle, or finding the position of shuffled image patches. Generative methods use autoencoders or generative adversarial networks (GANs). They are assessed by the quality of reconstructing the original images. Contrastive methods rely on using data augmentation to create positive pairs of an image. Different images are treated as negative pairs. Various distance metrics have been proposed to keep positive pairs close and negative pairs distant in the embedding space. Some examples include SimCLR [11], MoCo [32], BYOL [5], SimSiam [12], DINO [8]. Self-prediction methods mask out portions of the original image and try to reconstruct the original image. The main difference between self-prediction and generative methods is that self-prediction applies augmentation on portions of the original image while generative methods apply to the whole image. SSL application in medical imaging is rapidly rising, as medical labels are costly and time-consuming to obtain. SSL can speed up the process of model development.
II-B Unsupervised domain adaptation
@@
UDA methods can be divided into two types, moment matching and adversarial training. Moment matching methods try to match the first or second moments between the hidden activation distributions of the source and target domains [78]. Some examples are deep adaptation networks (DAN) [48], joint adaptation networks (JAN) [49], correlation alignment [66], and central moment discrepancy (CMD) [78]. Adversarial training methods try to learn domain invariant features. It trains a domain classifier to put the source and target domain features in the same space. The most used adversarial training method is domain adversarial neural networks (DANN) [25]. Other methods build upon it such as adversarial discriminative domain adaptation (ADDA) [70], maximum classifier discrepancy (MCD) [57], batch spectral penalization (BSP) [13], and minimum class confusion (MCC) [39].
@@
There are works of applying UDA in medical image analysis focusing on classification [1], object localization, and segmentation [41]. More specifically for skin lesion classification, Chamarthi et al. compare performance of different UDA methods [10]. Wang et al. demonstrate applying UDA methods can mitigate model bias against minority groups [73].
II-C Active domain adaptation
@@
UDA methods often exhibit lower performance compared to their supervised counterparts [67, 14]. Supervised domain adaptation (SDA) requires labels of the target domain. However, for the medical imaging field especially, obtaining large amounts of labels can be cost intensive. Therefore, selecting the most informative samples for annotation is crucial. This is where active learning (AL) comes in. AL is trying to maximize a model’s performance given a limited amount of labels. Active domain adaption (ADA) combines active learning with domain adaptation. Different active learning methods have different sampling strategies. Broadly, they are based on the two types of metrics, uncertainty [46, 59], and diversity [38, 35]. Uncertainty-based methods choose instances with the high uncertainty based on measures such as entropy, classification margins, or confidence [22, 24, 72]. Diversity-based methods choose instances that are representative of the entire dataset. They optimize for diversity in the embedding space with clustering or core-set selection [60, 26, 29, 62]. Several methods attempt to combine these two types of metrics [56, 36, 23, 2]. Zhao et al. applied SSL and AL for skin lesion segmentation, but there is no domain adaptation [79]. No prior work exists on applying ADA methods to skin lesion classification.
III Methodology
@@
Here, we propose a workflow combining SSL and DA methods to improve the generalization of skin lesion classification models. Figure 1 presents the steps of the workflow.
Figure 1: Proposed workflow.
@@
The initial step involves selecting an SSL pre-trained model on natural image datasets, such as ImageNet. Here we select the DINO SSL method since it’s shown to produce the SOTA performance on ImageNet among many SSL methods, e.g. SimCLR, MoCov2, Barlow Twins, BYOL, SwAV [50, 8]. Figure 2 shows the DINO architecture. It uses knowledge distillation with a student and teacher architecture. Aug 1 and Aug 2 represent different augmented views of an image, which are fed into the student and teacher networks, respectively. Both networks have the same model architecture. The loss term minimizes the cross-entropy between the features learned by the student and teacher networks. Centering and sharpening are applied on the teacher’s output to prevent the model from learning trivial solutions. It doesn’t require a large batch size or negative samples. We choose ResNet50 [33] as the backbone model because it’s used in other similar studies [10], allowing for direct comparison of results. The projector head has a couple linear layers with the hidden dimension as 256 and the output dimension as 65536. The DINO cross-entropy loss term between the teacher’s and student’s learned features is defined as follows:
@@
Figure 2: DINO SSL architecture.
The first step involves continuing DINO training on skin lesion datasets. Because no labels are required, training can be performed on all the available skin lesion data. This facilitates the model’s generalization from natural images to skin lesion data.
The second step involves finetuning the model on a labeled skin lesion dataset. Here, we focus on a common task: binary classification of melanoma (cancer) versus nevus (benign). We freeze the backbone model and fine-tune only the linear classification layer to isolate the impact of feature quality, excluding the influence of a more complex classifier.
The third step involves applying ADA. An optional preliminary step involves applying UDA to align the source and target domains without target domain labels. One popular UDA method is DANN [25], which has shown good performance among UDA methods applied to skin lesion datasets [10]. DANN is an adversarial training based method. It comprises three main components: a feature extractor, a class classifier, and a domain classifier. A gradient reversal layer between the domain classifier and the feature extractor encourages the feature extractor to learn domain invariant features.
Figure 3: Active domain adaptation architecture.
We compare several ADA methods: AADA [65], CLUE [56], BADGE [2], and random sampling. Figure 3 illustrates the general ADA architecture. We have skin lesion labels for the source domain dataset, but lack labels for the target domain datasets. Both source domain and target domain datasets can be input to the model, resulting in two computed losses. $L_{d}$ represents the domain loss, which we use to train the model to be unable to distinguish the domains. $L_{c}$ represents the skin lesion classification loss. Cross-entropy is used to calculate both losses. We apply active learning to select informative samples from the target domain, which are then sent to human experts for annotation. Using these annotated labels, we can then train the model using the classification loss. The domain classifier is optional; adversarial training-based ADA methods, such as AADA, utilize it, while other methods don’t.
For example, the active learning sampling criterion of AADA method is as such:
@@
It incorporates both the diversity cue $\frac{1-G_{d}(G_{f}(x))}{G_{d}(G_{f}(x))}$, and uncertainty cue $H(G_{y}(G_{f}(x)))$.
IV Experiment
To demonstrate the effectiveness of our approach in improving model generalization, we conduct tests on a set of skin lesion datasets exhibiting domain shifts. We compare our results with the existing approach using UDA methods. We demonstrate the benefits of SSL training and applying ADA respectively. When applying the ADA method, we consider real-world clinical usage by employing a small sampling size of 10. Given an additional labeling budget, we can employ an iterative training strategy with an increased number of provided samples.
IV-A Dataset
Fogelberg et al. [40] created skin lesion datasets representing different domains by grouping data from HAM10000 [69], BCN20000 [34], and MSK [9] according to biological traits such as patient age and lesion location. They quantified the domain shift intensity between groups using cosine similarity and Jensen-Shannon divergence. The details of the grouped datasets are presented in Table I. It lists a total of 11 datasets. Dataset H is the source domain, and the other 10 datasets are target domains. It also shows the number of samples in each dataset, the ratio of melanoma and nevus, and biological traits.
TABLE I: Skin lesion datasets of different domains.
Abbreviation Origin Biological factors Melanoma amount Nevus amount Total sample size
@@ IV-A Dataset
TABLE II: Comparison of our SSL methods together with the BSP UDA method and baseline reported from Chamarthi et al. [10].
HA HLH HLP B BA BLH BLP M MA MLH
Mel (ratio) 0.04 0.45 0.07 0.41 0.08 0.66 0.65 0.31 0.08 0.6
Baseline 0.14±0.02 0.69±0.04 0.37±0.15 0.57±0.02 0.19±0.06 0.73±0.03 0.77±0.05 0.34±0.01 0.15±0.04 0.68±0.03
BSP 0.16±0.03 0.82±0.02 0.65±0.04 0.75±0.01 0.34±0.05 0.86±0.01 0.83±0.02 0.46±0.02 0.17±0.03 0.73±0.01
Dino retrained 0.17±0.04 0.86±0.01 0.5±0.09 0.77±0.01 0.31±0.09 0.84±0.01 0.86±0.02 0.58±0.02 0.24±0.05 0.75±0.04
Dino pre-trained 0.13±0.04 0.89±0.03 0.6±0.15 0.76±0.01 0.34±0.07 0.83±0.01 0.83±0.02 0.52±0.01 0.12±0.02 0.77±0.01
SL pre-trained 0.16±0.04 0.79±0.03 0.71±0.12 0.73±0.01 0.33±0.05 0.83±0.02 0.80±0.02 0.52±0.01 0.20±0.02 0.70±0.02
IV-B Metrics
We use AUPRC metric because it effectively handels imbalanced datasets given greater weight to the cancer class. As shown in Table I, many of the skin lesion datasets are highly imbalanced, with the cancer class comprising as little as 4% of the data. We don’t use accuracy or AUROC because they are only suitable for balanced datasets. However, we must consider the varying AUPRC baselines across datasets, as these baselines reflect the different melanoma ratios.
IV-C Model Training